imqwerty (imqwerty):

2 equations 3variables l+m=1-n l^3 + m^3 = 1-n^2

rvc (rvc):

consider m=t where t is a parameter

OpenStudy (baru):

l=1 m=0 n=0

imqwerty (imqwerty):

it is one of the solutions (;

imqwerty (imqwerty):

rvc i didn't get it we put m=t then?

OpenStudy (baru):

every combo (1,0,0) (0,0,1) (0,1,0)

imqwerty (imqwerty):

there are more solutions :)

OpenStudy (jiteshmeghwal9):

\[l^2+m^2-lm=1+n\] this is the third eqn

OpenStudy (jiteshmeghwal9):

\[(l+m)(l^2+m^2-lm)=(1+n)(1-n)\]since\[l+m=1-n\]therefore\[l^2+m^2-lm=1+n\]

imqwerty (imqwerty):

we can clearly see that (0,0,1) is a solution But your equation is not satisfying this solution.

OpenStudy (jiteshmeghwal9):

well i have not done mistakes in using identities

OpenStudy (jiteshmeghwal9):

is (0,0,1) the answer given in ur book ?

imqwerty (imqwerty):

i will get the answers tomorrow my proff gave this prob to me . But according to my solution i am also not getting (0,01) but it is satisfying the equation

OpenStudy (jiteshmeghwal9):

i think if \[l+m=1+n\]then it is also satisfying my eqn because then i will get eqn \[l^2+m^2-lm=1-n\]

OpenStudy (jiteshmeghwal9):

\[l+m=1+n\]\[l^3+m^3=1-n^2\]\[\cancel{(l+m)}(l^2+m^2-lm)=\cancel {(1+n)}(1-n)\]

OpenStudy (jiteshmeghwal9):

the answer (0,0,1) is satisfying this eqn. actually my conclusion is that ur first eqn i.e l+m=1-n is probably wrong

imqwerty (imqwerty):

wait if (0,0,1) satisfies you equation but u cut that (l+m) with (1+n) l+m=0+0 and u can't divide by 0

OpenStudy (jiteshmeghwal9):

while i was cancelling i had assumed that i don't know the actual soln & cancelled them as variable.

OpenStudy (baru):

(1,-1,1) (-1,1,1)

imqwerty (imqwerty):

this will not satisfy the 2nd equation

OpenStudy (baru):

lhs and rhs =0 for both equations

imqwerty (imqwerty):

now wait its ok

OpenStudy (kainui):

do l,m,n have to only be positive integers or just integers?

imqwerty (imqwerty):

just integers

imqwerty (imqwerty):

i read the question wrong actually i gotta work on a new solution the question is correct tho

OpenStudy (baru):

they have to be integers? no fractions?

imqwerty (imqwerty):

yes only integers.

OpenStudy (kainui):

I'm getting stuck after factoring to this point which admittedly isn't very far lol: \[l^3+m^3=1-n^2\] \[(l+m)(l^2-lm+m^2)=(1-n)(1+n)\] \[l^2-lm+m^2=1+n\]

OpenStudy (baru):

there are solutions apart from the ones i mentioned?

imqwerty (imqwerty):

yes i know 6 solutions so far

OpenStudy (kainui):

If l and m are both even then n is odd, otherwise n is even from just looking at the equation in mod 2.

imqwerty (imqwerty):

yes :)

OpenStudy (baru):

\(n^2 =1-(l^3+m^3)\) l^3 and m^3 has to be a negative number implies L+m is negative and n=1-(L+m) n has to be positive

OpenStudy (kainui):

I made this just now to brute force check solutions like this with mods... heh... https://repl.it/B1t9

OpenStudy (kainui):

Well not solutions, it just tells you what n can be in mod k for every choice of l and m.

imqwerty (imqwerty):

try to eliminate n :)

OpenStudy (kainui):

\[l^2-lm+m^2+l+m-2=0\] It's an ellipse, so there are only finitely many points to check so I ask WA to do it for me cause it's so nice and it does and tells me: http://www.wolframalpha.com/input/?i=l%5E2%2Bm%5E2%2Bl%2Bm-lm-2%3D0&t=crmtb01&f=rc

OpenStudy (kainui):

Heh ok pretty bang-up job and not very elegant solution I'm giving here, but effectively done, maybe just have to check that it actually works... uhhh I'll stop here cause there's probably a prettier way and I don't wanna continue ruining it haha.

OpenStudy (baru):

well, now we have more than six solutions....

imqwerty (imqwerty):

but now (1,-1,1) and (-1,1,1) does not satisfy \[l^2-lm+m^2+l+m-2=0\]

imqwerty (imqwerty):

so the solutions reduce to 6 :D but (1,-1,1) and (-1,1,1) are solutions when we check them by putting in the original 2 equations

OpenStudy (dan815):

we have very few integers to check as its an ellipse

imqwerty (imqwerty):

we are not yet taught ellipse

OpenStudy (dan815):

are u taught about circle yet

imqwerty (imqwerty):

yes :)

OpenStudy (dan815):

ok then same thing xD

imqwerty (imqwerty):

x'D

imqwerty (imqwerty):

its algebra tho

OpenStudy (dan815):

ok ill think about a nicer solution

imqwerty (imqwerty):

:)

OpenStudy (dan815):

l+m=1-n l^3 + m^3 = 1-n^2 (l+m)(l^2+m^2-lm)=(1+n)(1-n) (l^2+m^2-lm)=(1+n) l^2+m^2=1+n+lm

OpenStudy (dan815):

brb

imqwerty (imqwerty):

ok :)

OpenStudy (anonymous):

what are those variables? Integers?

OpenStudy (anonymous):

Ah you've already mentioned it :)

OpenStudy (anonymous):

\(n=1\) gives \(l=-m\) as a family of solutions. For \(n \neq 1\) one can eliminate \(l\) easily to get\[(m+n-1)^2+m^2=1+n\]Simplify to get a quadratic in terms of \(m\):\[2m^2+(2n-2)m+n^2-3n=0\]look at the discriminant, in order to get integer values for \(m\) it must be non-negative and integer:\[\Delta_m=-n^2+4n+1 \ge0\]It gives \(n=2, 3, 4\) (why?) only, so you a few cases to check.

OpenStudy (anonymous):

I better be careful, last line must be: It gives \(n=0, 2, 3, 4\) (why?) only, so you have a few cases to check.

imqwerty (imqwerty):

yes we can conclude that n lies in the interval [0,4]

OpenStudy (baru):

one of the solutions kainui has shown is l=-3 m=-2 n=6

imqwerty (imqwerty):

ok ima tell the solution :)

imqwerty (imqwerty):

eliminating n from the equation we get-\[l^2+m^2+[1-(l+m)]^2=1\]after factorization-\[(l+m)(l^2 -lm+m^2+l+m-2)=0\]Now CASE 1---> suppose l+m=0 then n=1 and (l, m, n)=(-a, +a ,1) here a is an integer this will give us some solutions CASE 2----> suppose l+m≠0 then we must have -\[l^2 -lm +m^2+l+m-2=0\]and we ccan write it like this- :)\[(2l-m+1)^2+3(m+1)^2=12\]now we have 2 possibilities \[2l-m+1=0, y+1=\pm2\]and\[2l-m+1=\pm3,    y+1=\pm1\] Now analyzing all the case we get (l, m, n)=(1,0,0), (0,1,0), (-2,-3,6), (-3, -2, 6), (0, -2,3 ), (-2, 0, 3)

OpenStudy (pinklion23):

WOW